OpenStudy (anonymous):

. The formula to determine the _____ is based on the _____. (Points : 2) central limit theorem; standard deviation standard deviation; central limit theorem standard error; central limit theorem central limit theorem; z test

OpenStudy (anonymous):

Central limit theorem; standard deviation I believe ;o

ganeshie8 (ganeshie8):

trick question

ganeshie8 (ganeshie8):

idk the proof of central limit theorem yet, but CLT does use standard deviation

ganeshie8 (ganeshie8):

however standard error is a direct consequence of central limit theorem

ganeshie8 (ganeshie8):

CLT gives : standard error = \(\large Z^* \times \frac{s}{\sqrt{n}}\) right ?

ganeshie8 (ganeshie8):

option C is more fitting to above result...